An external force acting on a particle of mass m in one dimension is

given by F = βt2 where β is a positive constant. If the initial speed of the particle is

v1 (at t = 0) and its final speed is v2 at t = tf . Find constant β in terms of the given

quantities.


1
Expert's answer
2021-02-04T02:32:46-0500

Answer

Force

"F=m\frac{dv}{dt}={\beta t^2}"

"=\frac{dv}{dt}=\frac{\beta t^2}{m}"

By integration v limit v1 to v2 and time t =0 to t=tf

So we get

"v_2-v_1=\frac{\beta t^3 }{3m}"

So constant

"\beta=\frac{(v_2-v_1) 3m}{t_f^3}"